# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 5. Regularized solution
(5.1) $\begin{split}(\mathcal{R}^{*}g)(x)=&\int_{\mathcal{V}}g(y)W(x,y)\delta(p-\Phi(x,\alpha))\text{d}y,\ x\in\mathcal{U}.\end{split}$
As is well-known, assumption 3.2 ensures that $\mathcal{R}^{*}\mathcal{R}$ is an elliptic pseudo-differential operator $(\Psi$DO) [46, 19], [53, Section VIII.6.2] of order one in $\mathcal{U}$, i.e. $\mathcal{R}^{*}\mathcal{R}\in L^{1}(\mathcal{U})$.
@@
(5.2) $\begin{split}&\Psi(f+h)-\Psi(f)=\Psi^{\prime}(f;h)+O(\|h\|_{H^{1}(\mathcal{U})}^{2}),\ f,h\in H_{c}^{1}(\mathcal{U}_{b}),\\ &(1/2)\Psi^{\prime}(f;h)=(\mathcal{R}f-g_{\epsilon},\mathcal{R}h)_{L^{2}(\mathcal{V})}+\kappa\epsilon^{3}(\nabla f,\nabla h)_{L^{2}(\mathcal{U})},\end{split}$
and the functional $H_{c}^{1}(\mathcal{U}_{b})\ni h\to\Psi^{\prime}(f;h)$ is linear and continuous [41, p. 14]. Recall that $g_{\epsilon}$ is defined in (3.13). Then the subgradient of $\Psi$ coincides with the gradient [41, Proposition 3.20], and $f$ is the minimizer if and only if the gradient equals zero [41, Theorem 3.24]. The facts that $\Psi$ is proper and convex, which are required for the latter conclusion, are established in the proof of Lemma 3.5 in appendix A. Therefore the solution to (3.16) is the unique solution to the following equation, which is equivalent to the first order optimality condition:
(5.3) $(\mathcal{R}^{*}\mathcal{R}-\kappa\epsilon^{3}\Delta)f_{\epsilon}=\mathcal{R}^{*}g_{\epsilon},\ x\in\mathcal{U}_{b},\ f_{\epsilon}\in H_{c}^{1}(\mathcal{U}_{b}).$
This follows by setting $\Psi^{\prime}(f_{\epsilon};h)\equiv 0$, $h\in C_{0}^{\infty}(\mathcal{U}_{b})$, integrating by parts, and using that $C_{0}^{\infty}(\mathcal{U}_{b})$ is dense in $H_{c}^{1}(\mathcal{U}_{b})$.
@@
Since $\Psi(f\equiv 0)=\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}$ (cf. (3.16)), the solution $f_{\epsilon}$ satisfies $\|\mathcal{R}f_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq 2\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$ (cf. the paragraph above (3.16)). By construction, $f_{\epsilon}\in H_{c}^{1}(\mathcal{U}_{b})\subset H_{c}^{-1/2}(\mathcal{U}_{b})$, where $\mathcal{U}_{b}$ is bounded. Application of (3.6) completes the proof. ∎
Let $\tilde{D}$ be the complete symbol of $\mathcal{R}^{*}\mathcal{R}$, i.e. $\mathcal{R}^{*}\mathcal{R}=\text{Op}(\tilde{D}(x,\xi))$ (see [16, Theorem 3.4]). Rewrite (5.3) in an alternative form:
(5.5) $\begin{split}\big{(}\text{Op}(\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2})f_{\epsilon}\big{)}(x)=(\mathcal{R}^{*}g_{\epsilon})(x),\ x\in\mathcal{U}_{b}.\end{split}$
@@
(5.6) $\begin{split}&\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}f_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}g_{\epsilon}=:F_{\epsilon}.\end{split}$
@@
(5.7) $G_{\epsilon}:=\text{Op}\big{(}\big{[}\tilde{D}_{0}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon}.$
@@ 6.1. Preliminary calculations
(6.1) $\begin{split}F_{\epsilon}(x)&=(\mathcal{T}_{1}g_{\epsilon})(x),\ \mathcal{T}_{1}:=\text{Op}(\chi(x)|\xi|)\chi(z)\mathcal{R}^{*},\\ (\mathcal{T}_{1}g_{\epsilon})(x)&=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}|\xi|e^{i\xi\cdot(z-x)}\chi(z)\\ &\hskip 28.45274pt\times\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{V}}W(z,y)e^{i\nu(p-\Phi(z,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu\text{d}z\text{d}\xi.\end{split}$
Clearly, $\mathcal{T}_{1}\in I^{1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$. Here the canonical relation $\mathcal{C}^{*}$ is the image of $\mathcal{C}$ (see (3.3)) under the exchange of $T^{*}\mathcal{U}$ and $T^{*}\mathcal{V}$. Next, consider the expression
(6.2) $\begin{split}\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}\chi(z)W(z,y)|\xi|e^{i\xi\cdot(z-x)-i\nu(\Phi(z,\alpha)-\Phi(x,\alpha))}\text{d}z\text{d}\xi.\end{split}$
Upon changing variables $\xi\to\xi_{1}=\xi/\nu$, the stationary point of the phase satisfies $z=x$, $\xi_{1}=\Phi_{z}^{\prime}(z,\alpha)$, and the stationary point is nondegenerate. By the stationary phase method (see also [53, Ch. VI, eqs. (4.1)–(4.4)]),
(6.3) $\mathcal{T}_{1}=\mathcal{T}_{11}+\mathcal{T}_{12},\ \mathcal{T}_{11}\in I^{1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*}),\ \mathcal{T}_{12}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*}),$
where
(6.4) $\begin{split}&(\mathcal{T}_{11}g_{\epsilon})(x):=\frac{1}{2\pi}\int_{\mathbb{R}}|\nu|\int_{\mathcal{V}}K(x,\alpha)e^{i\nu(p-\Phi(x,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu,\ x\in\mathcal{U}_{0},\\ &K(x,\alpha):=\chi^{2}(x)|\Phi_{x}^{\prime}(x,\alpha)|W(x,(\Phi(x,\alpha),\alpha)),\end{split}$
and $\mathcal{T}_{12}:=\mathcal{T}_{1}-\mathcal{T}_{11}$. The inclusion $\mathcal{T}_{12}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$ follows by considering lower order terms in the stationary phase expansion as $\nu\to\infty$ of the integral in (6.2) [53, Chapter VIII, eqs. (2.14), (2.16)].
By the continuity of FIOs in Hölder-Zygmund spaces (see [23, Proposition 5.5] and (C.12), (C.13)), $\|\mathcal{T}_{12}g_{\epsilon}\|_{C^{1/2}(\mathcal{U})}<c$ for all $0<\epsilon\ll 1$. By Lemma 4.2, $\Delta g,\Delta g_{\epsilon}\in C_{0}^{3/2}(\mathcal{V})$ and $\|\Delta g_{\epsilon}\|_{C^{3/2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$. In this case $\|\mathcal{T}_{1}\Delta g_{\epsilon}\|_{C^{1/2}(\mathcal{U})}<c$, $0<\epsilon\ll 1$.
@@
(6.5) $\begin{split}F_{\epsilon}^{(l)}(x):=-\frac{1}{\pi}\int_{\mathcal{V}}&\frac{\chi_{1\text{D}}(p-\Phi(x,\alpha))K(x,\alpha)}{p-\Phi(x,\alpha)}\\ &\times\sum_{j}\partial_{p}\varphi_{\epsilon}(y-y_{j})a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}y.\end{split}$
Here and in the rest of this section, the pair $a_{0}(y),P(\alpha)$ stands for any one of the pairs $a_{l},P_{l}$ described in Lemma 4.2. This means that the domain of integration $\mathcal{V}$ in (6.5) can be replaced by a small neighborhood of $\text{supp}\,a_{l}$. This increase in support is due to interpolation.
To get (6.5) from (6.4) we expressed $\text{Op}(|\nu|)=-1/(\pi t^{2})$, integrated by parts with respect to $p$, and inserted a 1D cutoff function $\chi_{1\text{D}}$. The latter identically equals one in a neighborhood of zero. Also, we omitted smooth terms with uniformly bounded $\|\cdot\|_{C^{1}(\mathcal{U})}$ norm, $0<\epsilon\ll 1$, which arise due to the cutoff.
@@
for any fixed $A>0$.
@@ 6.2. Simplification of $F_{\epsilon}^{(l)}$, the leading order term of $F_{\epsilon}$
(6.7) $\begin{split}&Q(x,\alpha):=\Phi(x,\alpha)-P(\alpha).\end{split}$
Since we want to prove that $F_{\epsilon}^{(l)}(x)$ is uniformly bounded for all $x\in\mathcal{U}$, we drop $x$ from notation (e.g., $K(x,\alpha)$ of (6.4) becomes $K(\alpha)$) and make sure that all the estimates are uniform with respect to $x$. It suffices to assume that $x$ is confined to a sufficiently small open subset $U^{\prime}\subset\mathcal{U}$.
Change variables $p=w+\Phi(\alpha)$ in (6.5):
(6.8) $\begin{split}F_{\epsilon}^{(l)}=-\frac{1}{\pi}\sum_{j_{1}}\int_{\mathbb{R}}&\frac{\chi_{1\text{D}}(w)}{w}\sum_{j_{2}}\bigg{[}\partial_{w}\int_{\mathbb{R}}K(\alpha)\varphi_{\epsilon}(y-y_{j})\text{d}\alpha\bigg{]}\\ &\times a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}w,\ y=(\alpha,w+\Phi(\alpha)).\end{split}$
The sum with respect to $j_{1}$ is over $\alpha_{j_{1}}$ in the domain of the locally defined function $P(\alpha)$. Consider the integral with respect to $\alpha$:
(6.9) $\begin{split}\int_{\mathbb{R}}&K(\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\partial_{w}\varphi_{p}\bigg{(}\frac{w+\Phi(\alpha)-p_{j_{2}}}{\epsilon}\bigg{)}\text{d}\alpha\\ =&\epsilon^{-1}\int_{\mathbb{R}}K(\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\varphi_{p}^{\prime}\bigg{(}\frac{\Phi(\alpha)-\Phi(\alpha_{j_{1}})}{\epsilon}+u\bigg{)}\text{d}\alpha,\\ u:=&(w+\Phi(\alpha_{j_{1}})-p_{j_{2}})/\epsilon.\end{split}$
@@
(6.10) $\begin{split}h(u;\theta,\epsilon):=&\mu\int_{\mathbb{R}}K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)\varphi_{p}\bigg{(}\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}+u\bigg{)}\text{d}s.\end{split}$
@@
(6.11) $H(\check{t},\check{q};\theta,\epsilon):=\sum_{j_{2}}h^{\prime}(\check{t}+\check{q}-j_{2};\theta,\epsilon)a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
@@
(6.12) $\check{t}=(w+Q(\alpha_{j_{1}}))/\epsilon,\ \check{q}=P(\alpha_{j_{1}})/\epsilon,\ \theta=\alpha_{j_{1}}.$
@@
(6.13) $\begin{split}&\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(w)}{w}H\bigg{(}\frac{w+Q(\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};\theta,\epsilon\bigg{)}\text{d}w\\ &=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}H\bigg{(}\check{w}+\frac{Q(\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};\theta,\epsilon\bigg{)}\text{d}\check{w},\end{split}$
@@
(6.14) $\begin{split}G(\check{t},\check{q};\theta,\epsilon):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{H(\check{w}+\check{t},\check{q};\theta,\epsilon)}{\check{w}}\text{d}\check{w}.\end{split}$
@@
(6.15) $\begin{split}F_{\epsilon}^{(l)}=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}}G\bigg{(}\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon},\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
@@
Let $\text{dom}\,P$ denote the domain of the selected $P(\alpha)$. The following lemma is proven in appendix D.1.
@@
(6.16) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\begin{cases}O(\epsilon^{1/2})+O(\check{t}^{-1}),&\check{t}\to\infty,\\ O((-\check{t})^{-1/2}),&\check{t}\to-\infty.\end{cases}\end{split}$
In the above formula the big-$O$ term is uniform in $\check{q}\in\mathbb{R}$, $\theta\in\text{dom}\,P$, $|\check{t}|\leq A/\epsilon$, and $0<\epsilon\ll 1$.
@@
Combining Lemma 6.3 with (6.6) proves Lemma 6.1.

### v2

@@ 5. Regularized solution
(5.1) $\begin{split}(\mathcal{R}^{*}g)(x)=&\int_{\mathcal{V}}g(y)W(x,y)\delta(p-\Phi(x,\alpha))\text{d}y\\ =&\int_{\mathcal{I}_{\alpha}}g(\alpha,\Phi(x,\alpha))W(x,(\alpha,\Phi(x,\alpha)))\text{d}\alpha,\ x\in\mathcal{U}.\end{split}$
As is well-known, Assumption 3.2 ensures that $\mathcal{R}^{*}\mathcal{R}$ is an elliptic pseudo-differential operator $(\Psi$DO) [49, 22], [56, Section VIII.6.2] of order -1 in $\mathcal{U}$, i.e. $\mathcal{R}^{*}\mathcal{R}\in L^{-1}(\mathcal{U})$.
@@
(5.2) $\begin{split}&\Psi(f+h)-\Psi(f)=\Psi^{\prime}(f;h)+O(\|h\|_{H^{1}(\mathcal{U})}^{2}),\ f,h\in H_{0}^{1}(\mathcal{U}_{b}),\\ &(1/2)\Psi^{\prime}(f;h)=(\mathcal{R}f-\hat{f}_{\epsilon},\mathcal{R}h)_{L^{2}(\mathcal{V})}+\kappa\epsilon^{3}(\partial_{x}f,\partial_{x}h)_{L^{2}(\mathcal{U})},\end{split}$
and the functional $H_{0}^{1}(\mathcal{U}_{b})\ni h\to\Psi^{\prime}(f;h)$ is linear and continuous [45, p. 14]. Recall that $\hat{f}_{\epsilon}$ is defined in (3.14). Then the subgradient of $\Psi$ coincides with the gradient [45, Proposition 3.20], and $f$ is the minimizer if and only if the gradient equals zero [45, Theorem 3.24]. The fact that $\Psi$ is proper and convex, which is required for the latter conclusion, is established in the proof of Lemma 3.5 in Appendix A. Therefore the solution to (3.17) is the unique solution to the following equation, which is equivalent to the first order optimality condition:
(5.3) $(\mathcal{R}^{*}\mathcal{R}-\kappa\epsilon^{3}\Delta)f_{\epsilon}=\mathcal{R}^{*}\hat{f}_{\epsilon},\ x\in\mathcal{U}_{b},\ f_{\epsilon}\in H_{0}^{1}(\mathcal{U}_{b}).$
This follows by setting $\Psi^{\prime}(f_{\epsilon};h)\equiv 0$, $h\in C_{0}^{\infty}(\mathcal{U}_{b})$, integrating by parts, and using that $C_{0}^{\infty}(\mathcal{U}_{b})$ is dense in $H_{0}^{1}(\mathcal{U}_{b})$.
@@
Since $\Psi(f\equiv 0)=\|\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}$ (cf. (3.17)), the solution $f_{\epsilon}$ satisfies $\|\mathcal{R}f_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq 2\|\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$ (cf. the paragraph preceding (3.17)). For clarity, $\mathcal{R}f_{\epsilon}$ is the GRT of the solution $f_{\epsilon}$, while $\hat{f}_{\epsilon}$ is the interpolated version of the exact discrete data, i.e. $\hat{f}_{\epsilon}=(\mathcal{R}f)_{\epsilon}$. By construction, $f_{\epsilon}\in H_{0}^{1}(\mathcal{U}_{b})\subset H_{0}^{-1/2}(\mathcal{U}_{b})$, where $\mathcal{U}_{b}$ is bounded. Application of (3.6) completes the proof. ∎
Let $\tilde{D}$ be the complete symbol of $\mathcal{R}^{*}\mathcal{R}$, i.e. $\mathcal{R}^{*}\mathcal{R}=\text{Op}(\tilde{D}(x,\xi))$ (see [19, Theorem 3.4]). Rewrite (5.3) in an alternative form:
(5.5) $\begin{split}\big{(}\text{Op}(\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2})f_{\epsilon}\big{)}(x)=(\mathcal{R}^{*}\hat{f}_{\epsilon})(x),\ x\in\mathcal{U}_{b}.\end{split}$
@@
(5.6) $\begin{split}&\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}f_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}\hat{f}_{\epsilon}=:F_{\epsilon}.\end{split}$
@@
(5.7) $\mathcal{G}_{\epsilon}:=\text{Op}\big{(}\big{[}\tilde{D}_{0}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon}.$
@@ 6. Proving the boundedness of $F_{\epsilon}$
By the linearity of (5.6), to prove the lemma it suffices to consider $F_{\epsilon}$ as defined on the right in (5.6), but with $\hat{f}_{\epsilon}$ replaced by $g_{\epsilon}$ using a local piece of the data $g=\chi_{V}\mathcal{R}f$ introduced in Lemma 4.2.
@@ 6.1. Preliminary calculations
(6.1) $\begin{split}F_{\epsilon}(x)&=(\mathcal{T}g_{\epsilon})(x),\ \mathcal{T}:=\text{Op}(\chi(x)|\xi|)\chi(z)\mathcal{R}^{*},\\ (\mathcal{T}g_{\epsilon})(x)&=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}|\xi|e^{i\xi\cdot(z-x)}\chi(z)\\ &\hskip 28.45274pt\times\frac{1}{2\pi}\int_{\mathbb{R}}\int_{V}W(z,y)e^{i\nu(p-\Phi(z,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu\text{d}z\text{d}\xi.\end{split}$
Next, consider the integral
(6.2) $\begin{split}K_{1}(x,\alpha,\nu):=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}\chi(z)W(z,y)|\xi|e^{i\xi\cdot(z-x)-i\nu(\Phi(z,\alpha)-\Phi(x,\alpha))}\text{d}z\text{d}\xi.\end{split}$
Upon changing variables $\xi\to\xi_{1}=\xi/\nu$, the stationary point of the phase satisfies $z=x$, $\xi_{1}=\text{d}_{x}\Phi(x,\alpha)$, and the stationary point is nondegenerate. By the stationary phase method,
(6.3) $\begin{split}&K_{1}(x,\alpha,\nu)=|\nu|K(x,\alpha)+\Delta K(x,\alpha;\nu),\\ &K(x,\alpha):=\chi^{2}(x)|\text{d}_{x}\Phi(x,\alpha)|W(x,(\alpha,\Phi(x,\alpha))),\ x\in\mathcal{U},\\ &K\in C_{0}^{\infty}(\mathcal{U}_{b}\times\mathcal{I}_{\alpha}),\ \Delta K\in C_{0}^{\infty}(\mathcal{U}_{b}\times\mathcal{I}_{\alpha}\times(\mathbb{R}\setminus 0)),\ \Delta K\in S^{0}(\mathbb{R}).\end{split}$
By (6.1)–(6.3),
(6.4) $\begin{split}(\mathcal{T}g_{\epsilon})(x)=&\int_{I_{\alpha}}K(x,\alpha)(\Lambda g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha\\ &+\int_{I_{\alpha}}(\Delta\mathcal{K}_{x,\alpha}g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha,\ x\in\mathcal{U}.\end{split}$
Here $\Lambda\in L^{1}(\mathcal{V})$, and $\mathcal{K}_{x,\alpha}\in L^{1}(\mathcal{V})$ is a $\Psi$DO, which depends smoothly on $(x,\alpha)\in\mathcal{U}_{b}\times\mathcal{I}_{\alpha}$,
(6.5) $\begin{split}(\Lambda g)(\alpha,t)=&\frac{1}{2\pi}\int_{\mathcal{I}_{p}}|\nu|e^{i\nu(p-t)}g(\alpha,p)\text{d}p\text{d}\nu,\\ (\mathcal{K}_{x,\alpha^{\prime}}g)(\alpha,t)=&\frac{1}{2\pi}\int_{\mathcal{I}_{p}}\Delta K(x,\alpha^{\prime};\nu)e^{i\nu(p-t)}g(\alpha,p)\text{d}p\text{d}\nu,\ (\alpha,t)\in\mathcal{V}.\end{split}$
By Lemma 4.2,
[…]
Recall that $B_{\infty,\infty}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$ [1, Remark 6.4(2)]. By the continuity of $\Psi$DOs in Hölder-Zygmund spaces (see [1, Theorem 6.19]), $\|\Lambda\Delta g_{\epsilon}\|_{C^{1/2}(V_{1})}<c$, $\|\Delta\mathcal{K}_{x,\alpha^{\prime}}g_{\epsilon}\|_{C^{1/2}(V_{1})}<c$, $0<\epsilon\ll 1$, for any compact $V_{1}\subset\mathcal{V}$. It is now obvious that
(6.7) $\begin{split}\int_{I_{\alpha}}(\Lambda\Delta g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha+\int_{I_{\alpha}}(\Delta\mathcal{K}_{x,\alpha}g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha\in C_{0}^{1/2}(\mathcal{U}),\end{split}$
with its $C^{1/2}(\mathcal{U})$ norm uniformly bounded for $0<\epsilon\ll 1$.
@@
(6.8) $\begin{split}F_{\epsilon}^{(l)}(x):=&-\frac{1}{\pi}\int_{V}\frac{\chi_{1\text{D}}(p-\Phi(x,\alpha))K(x,\alpha)}{p-\Phi(x,\alpha)}\\ &\times\sum_{y_{j}\in V}\partial_{p}\varphi_{\epsilon}(y-y_{j})a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}y,\ x\in\mathcal{U}_{b}.\end{split}$
Recall that $K(x,\alpha)\equiv 0$ for any $x\in\mathcal{U}\setminus\mathcal{U}_{b}$, cf. (6.3). To get (6.8) from (6.4), (6.5), we used the identity $\text{Op}(|\nu|)=-1/(\pi t^{2})$, integrated by parts with respect to $p$, and inserted a 1D cutoff function $\chi_{1\text{D}}$. The latter identically equals one in a neighborhood of zero. Also, we omitted smooth terms with uniformly bounded $\|\cdot\|_{C^{1}(\mathcal{U})}$ norm, $0<\epsilon\ll 1$, which arise due to the cutoff.
@@
for any fixed $A>0$. The $O(\epsilon^{1/2})$ term is uniform in $\check{x}$.
@@ 6.2. Simplification of $F_{\epsilon}^{(l)}$, the leading order term of $F_{\epsilon}$
(6.10) $\begin{split}&Q(x,\alpha):=\Phi(x,\alpha)-P(\alpha)=\Phi(x,\alpha)-\Phi(x(\alpha),\alpha).\end{split}$
To prove that $F_{\epsilon}^{(l)}(x)$ is uniformly bounded for all $x\in\mathcal{U}_{b}$, it suffices to assume that $x$ is confined to a sufficiently small open subset $U^{\prime}\subset\mathcal{U}_{b}$.
Change variable $p=w+\Phi(\alpha)$ in (6.8):
(6.11) $\begin{split}F_{\epsilon}^{(l)}(x)=-\frac{1}{\pi}\sum_{j_{1}\in J_{\alpha}}\int_{\mathbb{R}}&\frac{\chi_{1\text{D}}(w)}{w}\sum_{j_{2}\in J_{p}}\bigg{[}\partial_{w}\int_{\mathcal{I}_{\alpha}}K(x,\alpha)\varphi_{\epsilon}(y-y_{j})\text{d}\alpha\bigg{]}\\ &\times a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}w,\ y=(\alpha,w+\Phi(\alpha)).\end{split}$
Consider the integral with respect to $\alpha$:
(6.12) $\begin{split}\int_{\mathcal{I}_{\alpha}}&K(x,\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\partial_{w}\varphi_{p}\bigg{(}\frac{w+\Phi(\alpha)-p_{j_{2}}}{\epsilon}\bigg{)}\text{d}\alpha\\ =&\epsilon^{-1}\int_{\mathcal{I}_{\alpha}}K(x,\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\varphi_{p}^{\prime}\bigg{(}\frac{\Phi(\alpha)-\Phi(\alpha_{j_{1}})}{\epsilon}+u\bigg{)}\text{d}\alpha,\\ u:=&(w+\Phi(\alpha_{j_{1}})-p_{j_{2}})/\epsilon.\end{split}$
@@
(6.13) $\begin{split}h(u;*):=&\mu\int_{\mathbb{R}}K(x,\theta+\epsilon\mu s)\varphi_{\alpha}(s)\varphi_{p}\bigg{(}\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}+u\bigg{)}\text{d}s,\\ *=&(x,\theta,\epsilon),\ u\in\mathbb{R},\ x\in U^{\prime},\ \theta\in I_{\alpha},\ 0<\epsilon\ll 1.\end{split}$
@@
(6.14) $H(\check{t},\check{q};*):=\sum_{j_{2}\in J_{p}}h^{\prime}(\check{t}+\check{q}-j_{2};*)a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
@@
(6.15) $\check{t}=(w+Q(x,\alpha_{j_{1}}))/\epsilon,\ \check{q}=P(\alpha_{j_{1}})/\epsilon,\ \theta=\alpha_{j_{1}},\ j_{1}\in J_{\alpha}.$
@@
(6.16) $\begin{split}&\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(w)}{w}H\bigg{(}\frac{w+Q(x,\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};*\bigg{)}\text{d}w\\ &=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}H\bigg{(}\check{w}+\frac{Q(x,\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};*\bigg{)}\text{d}\check{w},\end{split}$
@@
(6.17) $\begin{split}G(\check{t},\check{q};*):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{H(\check{w}+\check{t},\check{q};*)}{\check{w}}\text{d}\check{w}.\end{split}$
@@
(6.18) $\begin{split}F_{\epsilon}^{(l)}(x)=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}\in J_{\alpha}}G\bigg{(}\frac{Q(x,\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};x,\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
@@
The following lemma is proven in Appendix D.1.
@@
(6.19) $\begin{split}|G(\check{t},\check{q};*)|\leq&c\begin{cases}\epsilon^{1/2}+(1+\check{t})^{-1},&\check{t}>0,\\ (1-\check{t})^{-1/2},&\check{t}<0,\end{cases}\end{split}$
for some $c$ independent of $|\check{t}|\leq A/\epsilon$, $|\check{q}|\leq A/\epsilon$, $x\in U^{\prime}$, $\theta\in I_{\alpha}$, and $0<\epsilon\ll 1$.
@@
The number of segments $\Gamma_{l}$ described above Lemma 4.2 is uniformly bounded, so combining Lemma 6.3 with (6.9) proves Lemma 6.1.